OpenStudy (anonymous):

A can of white paint costs $16 while a can of green paint costs $20. Susan is mixing 12 cans of paint and will pay $19 per can for the mixture. How many cans of white paint will she purchase?

OpenStudy (anonymous):

Let g= green paint Let w= white paint So g+w=12 20g +16w= (12)(19) Rearrange the first equation g=12-w Substitute into the second equation. 20(12-w) +16w= 228 Solve for w 240 -20w +16w=228 12=4w 3=w So she bought 3 cans of white paint and 9 cans of green.